# Accelerating Training of Recursive Reasoning Models with Curriculum Guided Adaptive Recursion

### v2

@@
Abstract
@@
I Introduction
The recent surge in large language models (LLMs) with hundreds of billions of parameters has demonstrated remarkable capabilities across diverse tasks [1, 2]. However, this paradigm of scaling model size incurs prohibitive computational costs during both training and inference, limiting accessibility and deployability. An alternative direction has emerged through test-time computation scaling [3, 4], where smaller models iteratively refine their outputs through multiple reasoning steps, trading inference cycles for model parameters.
Building upon this principle, the recently proposed Hierarchical Reasoning Model (HRM) [5] and its simplified variant, Tiny Recursive Model (TRM) [6], have shown that networks with merely 7M parameters can match or exceed LLMs ten thousand times their size on hard reasoning tasks such as Sudoku, maze solving, and ARC-AGI [7]. The key insight is recursive reasoning: a tiny network iteratively refines solutions through nested recursion cycles and deep supervision [8], effectively emulating deep architectures while maintaining parameter efficiency. Through adaptive computation time [9], TRM achieves strong performance on reasoning benchmarks using only 7M parameters. However, training these models remains computationally expensive, with the original TRM paper reporting approximately 36 GPU-hours for Sudoku extreme dataset, limiting broader adoption and rapid experimentation.
@@ I Introduction
We propose CGAR (Curriculum-Guided Adaptive Recursion), a novel training methodology that fundamentally rethinks how recursive models learn. Unlike all prior curriculum learning approaches that focus on data ordering [10, 11, 12], sample weighting, or parameter reduction [13], CGAR introduces the first curriculum on architectural recursion depth itself. This paradigm shift recognizes that the effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ of recursive architectures is a learnable training hyperparameter that should adapt with optimization progress, rather than remain fixed throughout training. CGAR introduces two synergistic innovations. First, Progressive Depth Curriculum (PDC) dynamically schedules recursion parameters $(n,T)$ based on normalized training progress $\rho=e/E$, implementing a three-stage curriculum that transitions from shallow $(2,1)$ through medium $(4,2)$ to full depth $(6,3)$ configurations, providing 41.4% FLOPs reduction while preventing early-stage overfitting. Second, Hierarchical Supervision Weighting (HSW) applies principled exponential decay $w_{t}=\lambda^{t-1}/Z_{\lambda}$ to supervision steps, derived from empirical observations of gradient magnitude decay in recursive architectures, achieving 40% gradient variance reduction and accelerated convergence through improved signal-to-noise ratio in stochastic gradients.
@@
The remainder of this paper is organized as follows: Section II reviews related work. Section III presents the CGAR framework. Section IV describes experimental setup. Section V reports results and ablations. Section VI concludes with limitations and future directions. The source code and trained models will be made publicly available upon publication.
Figure 1: Illustration of CGAR architecture follows TRM with recursive transformer blocks. CGAR introduces two key modifications: Progressive Depth Curriculum (PDC) adjusts recursion depth $(n,T)$ across training phases, and Hierarchical Supervision Weighting (HSW) applies exponential decay $w_{t}=\lambda^{t-1}$ to supervision losses.
II Related Work
@@
Curriculum learning [10] trains neural networks on progressively difficult examples, improving convergence and generalization across diverse domains [14, 12, 15]. Progressive architectures adjust network capacity during training: Rusu et al. [16] incrementally expand capacity for continual learning, while Huang et al. [17] introduced stochastic depth to vary effective network depth. Recent work applies curriculum to neural architecture search [18, 19]. Tang et al. [20] investigated progressive learning from shallow to deep representations via layer-wise pretraining. Training efficiency has been addressed through architectural innovations [21, 13, 22, 23]. Cao and Tsang [24] proposed adaptive momentum for training, eliminating hyperparameter tuning. However, existing curriculum methods focus on data ordering, sample weighting, or parameter reduction, treating architectural depth as fixed.
@@
Adaptive computation mechanisms dynamically allocate resources based on input complexity. Graves [9] introduced Adaptive Computation Time (ACT) for RNNs with learned halting at inference. PonderNet [25] learns variable computation steps, while early exit strategies reduce inference costs in CNNs [26] and transformers [27]. Raposo et al. [28] proposed Mixture-of-Depths for language models. Han et al. [29] survey dynamic networks adjusting depth, width, or resolution at inference, while Snell et al. [3] demonstrate test-time computation scaling for reasoning. Deep supervision [8] attaches auxiliary losses to intermediate layers for effective gradient flow. Chen et al. [30] introduced GradNorm for dynamic loss balancing, and Zhang et al. [31] investigated hierarchical supervision for multi-task learning. These approaches do not account for temporal decay in recursive architectures.
@@
Recursive reasoning mechanisms have gained attention for complex problem solving. Universal Transformers [32] apply recurrent processing to transformer blocks. Self-refinement methods [33, 34] enable iterative improvement through self-feedback. Sun et al. [35] explored recursive spawning, while Qasim et al. [36] developed recursive decomposition for enhanced LLM reasoning. Chen et al. [37] introduced Neural ODEs for continuous-depth networks, and Wang et al. [38] proposed Loop-Residual Networks. Most relevant, Chu et al. [5] developed the Hierarchical Reasoning Model (HRM) requiring approximately 72 GPU-hours training. Jolicoeur-Martineau [6] simplified HRM into TRM, demonstrating 7M parameter models matching LLMs $10{,}000\times$ larger, though training requires approximately 36 GPU-hours.
CGAR differs fundamentally from prior work. Unlike data-based curricula [10, 11, 12] and parameter-efficient methods [13], we introduce curriculum learning on architectural recursion depth $(n,T)$, dynamically adjusting effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ during training. Unlike inference-time adaptation [9, 25], our progressive depth curriculum operates at training time via deterministic scheduling. Unlike uniform supervision [6] and task-based weighting [31], our hierarchical weighting $w_{t}=\lambda^{t-1}$ accounts for gradient magnitude decay specific to recursive architectures. CGAR achieves $1.71\times$ training speedup while maintaining competitive accuracy, with ablations revealing progressive depth curriculum alone provides $2.26\times$ speedup, demonstrating benefits beyond computational savings.
III Method
@@
Our goal is to accelerate the training speed of Tiny Recursive Model (TRM) [6], which suffers from training inefficiency due to fixed-depth training at all epochs (the original TRM paper reports $\sim\!36$ GPU-hours per dataset on their hardware). We propose CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that achieves $1.71\times$ speedup through two complementary techniques: progressive depth curriculum (PDC) and hierarchical supervision weighting (HSW). Under controlled conditions on A100 GPU, CGAR reduces training time from $10.93$ hours (our replicated TRM baseline) to $6.38$ hours. We begin by reviewing TRM’s architecture and identifying its training inefficiencies, then present our solutions with mathematical formulations and theoretical analysis.
III-A Background: TRM Architecture and Training
@@
TRM operates on supervised learning tasks with dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i}^{*})\}_{i=1}^{N}$, where inputs $\bm{x}_{i}\in\mathbb{R}^{L\times V}$ have length $L$ over vocabulary size $V$, and outputs $\bm{y}_{i}^{*}\in\mathbb{R}^{L\times V}$. The model learns a function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ parameterized by $\theta\in\Theta\subset\mathbb{R}^{p}$ that minimizes expected cross-entropy loss $\mathbb{E}_{(\bm{x},\bm{y}^{*})\sim\mathcal{D}}[\ell_{\text{CE}}(f_{\theta}(\bm{x}),\bm{y}^{*})]$.
@@
$\bm{z}^{(t,j,k)}$ $=\mathcal{T}_{\theta}\left(\bm{x}_{\text{emb}}\oplus\bm{y}^{(t,j,k-1)}\oplus\bm{z}^{(t,j,k-1)}\right)$ (1)
$\bm{y}^{(t,j)}$ $=\mathcal{T}_{\theta}\left(\bm{y}^{(t,j-1)}\oplus\bm{z}^{(t,j,n)}\right)$ (2)
@@
TRM trains through deep supervision [8], attaching losses to all $N_{\text{sup}}$ steps with uniform weighting $w_{t}=1/N_{\text{sup}}$. Let $h_{\text{out}}:\mathbb{R}^{L\times D}\rightarrow\mathbb{R}^{L\times V}$ denote the output projection head. The training objective is:
$\mathcal{L}_{\text{TRM}}(\theta)=\frac{1}{N_{\text{sup}}}\sum_{t=1}^{N_{\text{sup}}}\ell_{\text{CE}}\left(h_{\text{out}}(\bm{y}^{(t,T,n)}),\bm{y}^{*}\right)$ (3)
@@
Problem I: Fixed-Depth Training Inefficiency
Figure 2: CGAR Training with Progressive Curriculum and Hierarchical Weighting
@@
Standard TRM training applies fixed recursion parameters $(\bar{n},\bar{T})=(6,3)$ uniformly across all epochs $e\in[E]$ and samples $(\bm{x},\bm{y}^{*})\in\mathcal{D}$, incurring total computational cost $\mathcal{C}_{\text{total}}=\mathcal{O}(EBLD^{2}\cdot\mathcal{D}_{\text{eff}}(\bar{n},\bar{T}))$ FLOPs where $B$ is batch size. This fixed-depth strategy wastes computation in two ways. First, during early training when parameters $\theta^{(e)}$ lie far from optimum $\theta^{*}\in\arg\min_{\theta}\mathbb{E}[\mathcal{L}(\theta)]$, the deep architecture with $\mathcal{D}_{\text{eff}}=42$ layers causes overfitting. Defining the generalization gap as $\mathcal{R}(\theta):=\mathbb{E}_{\mathcal{D}_{\text{test}}}[\ell(\theta)]-\mathbb{E}_{\mathcal{D}_{\text{train}}}[\ell(\theta)]$, we empirically observe $\mathcal{R}(\theta^{(e)})\propto\mathcal{D}_{\text{eff}}$ for early epochs $e<0.3E$. Second, samples vary in difficulty: if $\delta(\bm{x}):=\min\{t:h_{\text{out}}(\bm{y}^{(t)})=\bm{y}^{*}\}$ denotes minimum steps for correctness, then samples with $\delta(\bm{x})\ll N_{\text{sup}}$ waste computation on redundant refinement. On Sudoku-Extreme, $\mathbb{E}[\delta(\bm{x})]\approx 3.8$ while $N_{\text{sup}}=16$, suggesting $\sim\!76\%$ wasted steps.
Solution I: Progressive Depth Curriculum
To address fixed-depth inefficiency, we introduce Progressive Depth Curriculum (PDC), which adapts recursion parameters $(n,T)$ as training progresses. Define normalized progress $\rho:=e/E\in[0,1]$ for epoch $e\in[E]$. The curriculum function $\mathcal{C}_{\text{PDC}}:[0,1]\rightarrow\mathbb{N}^{2}$ maps progress to depth parameters via a piecewise-constant schedule with $K$ stages, transition thresholds $0=\tau_{0}<\tau_{1}<\cdots<\tau_{K}=1$, and stage-specific depths $(n_{i},T_{i})\in\mathbb{N}^{2}$ satisfying monotonicity $\mathcal{D}_{\text{eff}}(n_{1},T_{1})<\cdots<\mathcal{D}_{\text{eff}}(n_{K},T_{K})$:
$\mathcal{C}_{\text{PDC}}(\rho):=\sum_{i=1}^{K}(n_{i},T_{i})\cdot\mathds{1}_{[\tau_{i-1},\tau_{i})}(\rho)$ (4)
where $\mathds{1}_{A}(\cdot)$ is the indicator on set $A$. We instantiate a three-stage curriculum ($K=3$) with thresholds $(\tau_{1},\tau_{2})=(0.3,0.6)$ selected via validation grid search. The schedule starts with shallow recursion $(n_{1},T_{1})=(2,1)$ giving $\mathcal{D}_{\text{eff}}=6$ layers for $\rho\in[0,0.3)$, progresses to medium depth $(n_{2},T_{2})=(4,2)$ with $\mathcal{D}_{\text{eff}}=20$ for $\rho\in[0.3,0.6)$, and reaches full depth $(n_{3},T_{3})=(6,3)$ with $\mathcal{D}_{\text{eff}}=42$ for $\rho\in[0.6,1]$. This gradual deepening prevents early overfitting while enabling complex reasoning in later training.

### v3

# Accelerating Training Speed of Tiny Recursive Models with Curriculum Guided Adaptive Recursion

@@
Abstract.
Background: Recursive reasoning models achieve strong performance on complex reasoning tasks through iterative refinement, enabling tiny networks to match large language models thousands of times their size. However, training these networks remains computationally expensive, with prior work reporting 36 GPU-hours for Sudoku extreme, limiting broader adoption and research. Existing models employ fixed recursion depth uniformly across all training epochs and uniform supervision weighting across all reasoning steps, leading to inefficient training.
Objectives: We propose CGAR (Curriculum-Guided Adaptive Recursion), a novel training methodology that applies curriculum learning to architectural depth rather than traditional data ordering. CGAR introduces two synergistic components: Progressive Depth Curriculum (PDC) dynamically adjusts recursion depth from shallow to deep configurations during training and Hierarchical Supervision Weighting (HSW) applies exponentially decaying importance to supervision steps, aligning loss weighting with observed gradient magnitude decay.
Methods: Progressive Depth Curriculum implements a three-stage schedule that transitions from shallow $(2,1)$ through medium $(4,2)$ to full depth $(6,3)$ configurations based on normalized training progress, providing 41.4% FLOPs reduction while preventing early-stage overfitting. Hierarchical Supervision Weighting applies principled exponential decay $w_{t}=\lambda^{t-1}/Z_{\lambda}$ to supervision steps, achieving 40% gradient variance reduction and accelerated convergence through improved signal-to-noise ratio in stochastic gradients.
@@ Abstract.
Conclusions: CGAR demonstrates that principled curriculum on architectural depth enables efficient training of recursive reasoning models on modest hardware. By treating architectural depth as a curriculum-scheduled training parameter rather than a fixed constant, CGAR achieves substantial computational savings while preventing early-stage overfitting, making recursive reasoning models more practical for broader adoption in neurosymbolic AI, program synthesis and interpretable reasoning systems. Code and models are available at https://github.com/Kaleemullahqasim/CGAR and https://huggingface.co/Kaleemullah/trm-cgar-sudoku.
copyright: ccdoi: 10.1613/jair.1.xxxxxjournalvolume: 83article: 27publicationmonth: 8journalyear: 2025
1. Introduction
The recent surge in large language models (LLMs) with hundreds of billions of parameters has demonstrated strong capabilities across diverse tasks (undefb, ; undefg, ). However, this approach of scaling model size incurs prohibitive computational costs during both training and inference, limiting accessibility and deployability. An alternative direction has emerged through test-time computation scaling (undefab, ; undefq, ), where smaller models iteratively refine their outputs through multiple reasoning steps, trading inference cycles for model parameters.
Building upon this principle, the recently proposed Hierarchical Reasoning Model (HRM) (undefaf, ) and its simplified variant, Tiny Recursive Model (TRM) (undefp, ), have shown that networks with merely 7M parameters can match or exceed LLMs ten thousand times their size on hard reasoning tasks such as Sudoku, maze solving and ARC-AGI (undeff, ). The key insight is recursive reasoning: a tiny network iteratively refines solutions through nested recursion cycles and deep supervision (undefr, ), effectively emulating deep architectures while maintaining parameter efficiency. Through adaptive computation time (undefi, ), TRM achieves strong performance on reasoning benchmarks using only 7M parameters. However, training these models remains computationally expensive, with the original TRM paper reporting approximately 36 GPU-hours for Sudoku extreme dataset, limiting broader adoption and rapid experimentation.
@@ 1. Introduction
We propose CGAR (Curriculum-Guided Adaptive Recursion), a novel training methodology that fundamentally rethinks how recursive models learn. Unlike all prior curriculum learning approaches that focus on data ordering (undefa, ; undefl, ; undefac, ), sample weighting, or parameter reduction (undefn, ), CGAR introduces the first curriculum on architectural recursion depth itself. This approach recognizes that the effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ of recursive architectures is a learnable training hyperparameter that should adapt with optimization progress, rather than remain fixed throughout training. CGAR introduces two synergistic contributions. First, Progressive Depth Curriculum (PDC) dynamically schedules recursion parameters $(n,T)$ based on normalized training progress $\rho=e/E$, implementing a three-stage curriculum that transitions from shallow $(2,1)$ through medium $(4,2)$ to full depth $(6,3)$ configurations, providing 41.4% FLOPs reduction while preventing early-stage overfitting. Second, Hierarchical Supervision Weighting (HSW) applies principled exponential decay $w_{t}=\lambda^{t-1}/Z_{\lambda}$ to supervision steps, derived from empirical observations of gradient magnitude decay in recursive architectures, achieving 40% gradient variance reduction and accelerated convergence through improved signal-to-noise ratio in stochastic gradients.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work. Section 3 presents the CGAR framework. Section 4 describes experimental setup. Section 5 reports results and ablations. Section 6 concludes with limitations and future directions.
2. Related Work
@@
Curriculum learning (undefa, ) trains neural networks on progressively difficult examples, improving convergence and generalization across diverse domains (undefj, ; undefac, ; zhou2023curriculum). Progressive architectures adjust network capacity during training: Rusu et al. (undefy, ) incrementally expand capacity for continual learning, while Huang et al. (undefo, ) introduced stochastic depth to vary effective network depth. Recent work applies curriculum to neural architecture search (undefaj, ; undefk, ). Tang et al. (tang2023progressive) investigated progressive learning from shallow to deep representations via layer-wise pretraining. Training efficiency has been addressed through architectural advances (undefad, ; undefn, ; undefak, ; undefae, ). Cao and Tsang (undefc, ) proposed adaptive momentum for training, eliminating hyperparameter tuning. However, existing curriculum methods focus on data ordering, sample weighting, or parameter reduction, treating architectural depth as fixed.
@@
Adaptive computation mechanisms dynamically allocate resources based on input complexity. Graves (undefi, ) introduced Adaptive Computation Time (ACT) for RNNs with learned halting at inference. PonderNet (undef, ) learns variable computation steps, while early exit strategies reduce inference costs in CNNs (undefag, ) and transformers (undefah, ). Raposo et al. (undefx, ) proposed Mixture-of-Depths for language models. Han et al. (undefm, ) survey dynamic networks adjusting depth, width, or resolution at inference, while Snell et al. (undefab, ) demonstrate test-time computation scaling for reasoning. Deep supervision (undefr, ) attaches auxiliary losses to intermediate layers for effective gradient flow. Chen et al. (undefe, ) introduced GradNorm for dynamic loss balancing and Xu et al. (undefai, ) proposed multi-task guided prediction with intermediate auxiliary supervision. These approaches do not account for temporal decay in recursive architectures.
@@
Recursive reasoning mechanisms have gained attention for complex problem solving. Universal Transformers (undefh, ) apply recurrent processing to transformer blocks. Self-refinement methods (undefu, ; undefaa, ) enable iterative improvement through self-feedback. Sun et al. (undefz, ) explored recursive spawning, while Qasim et al. (undefw, ) developed recursive decomposition for improved LLM reasoning. Chen et al. (undefd, ) introduced Neural ODEs for continuous-depth networks and Wang et al. (undefv, ) proposed Loop-Residual Networks. Most relevant, Chu et al. (undefaf, ) developed the Hierarchical Reasoning Model (HRM) requiring approximately 72 GPU-hours training. Jolicoeur-Martineau (undefp, ) simplified HRM into TRM, demonstrating 7M parameter models matching LLMs $10{,}000\times$ larger, though training requires approximately 36 GPU-hours. CGAR differs fundamentally from prior work. Unlike data-based curricula (undefa, ; undefl, ; undefac, ) and parameter-efficient methods (undefn, ), we introduce curriculum learning on architectural recursion depth $(n,T)$, dynamically adjusting effective computational depth $\mathcal{D}_{\text{eff}}(n,T)$ during training. Unlike inference-time adaptation (undefi, ; undef, ), our progressive depth curriculum operates at training time via deterministic scheduling. Unlike uniform supervision (undefp, ), our hierarchical weighting $w_{t}=\lambda^{t-1}$ accounts for gradient magnitude decay specific to recursive architectures. CGAR achieves $1.71\times$ training speedup while maintaining competitive accuracy, with ablations revealing progressive depth curriculum alone provides $2.26\times$ speedup, demonstrating benefits beyond computational savings.
Figure 1. Illustration of CGAR architecture follows TRM with recursive transformer blocks. CGAR introduces two key modifications: Progressive Depth Curriculum (PDC) adjusts recursion depth $(n,T)$ across training phases and Hierarchical Supervision Weighting (HSW) applies exponential decay $w_{t}=\lambda^{t-1}$ to supervision losses.
3. Method
@@
We accelerate the training of Tiny Recursive Model (TRM) (undefp, ), which suffers from training inefficiency due to fixed-depth training at all epochs (the original TRM paper reports $\sim\!36$ GPU-hours per dataset on their hardware). We propose CGAR (Curriculum-Guided Adaptive Recursion), a training methodology that achieves $1.71\times$ speedup through two complementary techniques: progressive depth curriculum (PDC) and hierarchical supervision weighting (HSW). Under controlled conditions on A100 GPU, CGAR reduces training time from $10.93$ hours (our replicated TRM baseline) to $6.38$ hours. We begin by reviewing TRM’s architecture and identifying its training inefficiencies then present our solutions with mathematical formulations and theoretical analysis.
3.1. Background: TRM Architecture and Training
@@
TRM operates on supervised learning tasks with dataset $\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i}^{*})\}_{i=1}^{N}$, where inputs $\bm{x}_{i}\in\mathbb{R}^{L\times V}$ have length $L$ over vocabulary size $V$ and outputs $\bm{y}_{i}^{*}\in\mathbb{R}^{L\times V}$. The model learns a function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ parameterized by $\theta\in\Theta\subset\mathbb{R}^{p}$ that minimizes expected cross-entropy loss $\mathbb{E}_{(\bm{x},\bm{y}^{*})\sim\mathcal{D}}[\ell_{\text{CE}}(f_{\theta}(\bm{x}),\bm{y}^{*})]$.
@@
(1) $\bm{z}^{(t,j,k)}$ $=\mathcal{T}_{\theta}\left(\bm{x}_{\text{emb}}\oplus\bm{y}^{(t,j,k-1)}\oplus\bm{z}^{(t,j,k-1)}\right)$
(2) $\bm{y}^{(t,j)}$ $=\mathcal{T}_{\theta}\left(\bm{y}^{(t,j-1)}\oplus\bm{z}^{(t,j,n)}\right)$
@@
TRM trains through deep supervision (undefr, ), attaching losses to all $N_{\text{sup}}$ steps with uniform weighting $w_{t}=1/N_{\text{sup}}$. Let $h_{\text{out}}:\mathbb{R}^{L\times D}\rightarrow\mathbb{R}^{L\times V}$ denote the output projection head. The training objective is:
(3) $\mathcal{L}_{\text{TRM}}(\theta)=\frac{1}{N_{\text{sup}}}\sum_{t=1}^{N_{\text{sup}}}\ell_{\text{CE}}\left(h_{\text{out}}(\bm{y}^{(t,T,n)}),\bm{y}^{*}\right)$
@@
Figure 2. CGAR Training with Progressive Curriculum and Hierarchical Weighting
3.2. Problem I: Fixed-Depth Training Inefficiency
Standard TRM training applies fixed recursion parameters $(\bar{n},\bar{T})=(6,3)$ uniformly across all epochs $e\in[E]$ and samples $(\bm{x},\bm{y}^{*})\in\mathcal{D}$, incurring total computational cost $\mathcal{C}_{\text{total}}=\mathcal{O}(EBLD^{2}\cdot\mathcal{D}_{\text{eff}}(\bar{n},\bar{T}))$ FLOPs where $B$ is batch size. This fixed-depth strategy wastes computation in two ways. First, during early training when parameters $\theta^{(e)}$ lie far from optimum $\theta^{*}\in\arg\min_{\theta}\mathbb{E}[\mathcal{L}(\theta)]$, the deep architecture with $\mathcal{D}_{\text{eff}}=42$ layers causes overfitting. Defining the generalization gap as $\mathcal{R}(\theta):=\mathbb{E}_{\mathcal{D}_{\text{test}}}[\ell(\theta)]-\mathbb{E}_{\mathcal{D}_{\text{train}}}[\ell(\theta)]$, we empirically observe $\mathcal{R}(\theta^{(e)})\propto\mathcal{D}_{\text{eff}}$ for early epochs $e<0.3E$. Second, samples vary in difficulty: if $\delta(\bm{x}):=\min\{t:h_{\text{out}}(\bm{y}^{(t)})=\bm{y}^{*}\}$ denotes minimum steps for correctness then samples with $\delta(\bm{x})\ll N_{\text{sup}}$ waste computation on redundant refinement. On Sudoku-Extreme, $\mathbb{E}[\delta(\bm{x})]\approx 3.8$ while $N_{\text{sup}}=16$, suggesting $\sim\!76\%$ wasted steps.
3.3. Solution I: Progressive Depth Curriculum
To address fixed-depth inefficiency, we introduce Progressive Depth Curriculum (PDC), which adapts recursion parameters $(n,T)$ as training progresses. Define normalized progress $\rho:=e/E\in[0,1]$ for epoch $e\in[E]$. The curriculum function $\mathcal{C}_{\text{PDC}}:[0,1]\rightarrow\mathbb{N}^{2}$ maps progress to depth parameters via a piecewise-constant schedule with $K$ stages, transition thresholds $0=\tau_{0}<\tau_{1}<\cdots<\tau_{K}=1$ and stage-specific depths $(n_{i},T_{i})\in\mathbb{N}^{2}$ satisfying monotonicity $\mathcal{D}_{\text{eff}}(n_{1},T_{1})<\cdots<\mathcal{D}_{\text{eff}}(n_{K},T_{K})$:
(4) $\mathcal{C}_{\text{PDC}}(\rho):=\sum_{i=1}^{K}(n_{i},T_{i})\cdot\mathds{1}_{[\tau_{i-1},\tau_{i})}(\rho)$
where $\mathds{1}_{A}(\cdot)$ is the indicator on set $A$. We instantiate a three-stage curriculum ($K=3$) with thresholds $(\tau_{1},\tau_{2})=(0.3,0.6)$ selected via validation grid search. The schedule starts with shallow recursion $(n_{1},T_{1})=(2,1)$ giving $\mathcal{D}_{\text{eff}}=6$ layers for $\rho\in[0,0.3)$, progresses to medium depth $(n_{2},T_{2})=(4,2)$ with $\mathcal{D}_{\text{eff}}=20$ for $\rho\in[0.3,0.6)$ and reaches full depth $(n_{3},T_{3})=(6,3)$ with $\mathcal{D}_{\text{eff}}=42$ for $\rho\in[0.6,1]$. This gradual deepening prevents early overfitting while enabling complex reasoning in later training.